# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v1

@@ 5.2. A preliminary lemma on point counting
In what follows, if $x$ is a complex number, we use $\|x\|$ to denote its absolute value, while if $X$ is a set, we use $|X|$ to denote its cardinality. For $\mathscr{X}$ a Deligne-Mumford stack, we follow the standard convention that $|\mathscr{X}(\mathbb{F}_{q})|$ denotes $\sum_{x\in\mathscr{X}(\mathbb{F}_{q})}\frac{1}{|\operatorname{Aut}(x)|}$, so points are weighted inversely proportionally to their automorphism groups. Of course, when $\mathscr{X}$ is a scheme, this agrees with the usual notion of $|\mathscr{X}(\mathbb{F}_{q})|$.
@@
Fix a prime power $q$ and a sequence $\{Y_{n}\}_{n\in S}$ of nonempty smooth Deligne-Mumford stacks over $\operatorname{Spec}\mathbb{F}_{q}$ with $Y_{n}$ of pure dimension $n$, where $n$ traverses over an infinite sequence of increasing integers $S\subset\mathbb{Z}_{\geq 0}$. Suppose that the following cohomological conditions are satisfied:
@@ 5.3. Applying the lemma to count points on Hurwitz spaces
Suppose $G$ is a finite group, $c\subset G$ is a conjugacy class, and $(G,c)$ is non-splitting. Let $q$ be an prime power satisfying $\gcd(q,|G|)=1$. If $Y_{n}\subset\operatorname{CHur}^{G,c}_{n,\mathbb{F}_{q}}$ is a union of connected components which has $r$ geometrically irreducible components for each $n$ sufficiently large, there exist constants $C,I,J$, all independent of $n$ and $q$, but possibly depending on $r$, so that, as long as $q>C^{2}$,
@@
once we exhibit a suitable normal crossings compactification of configuration space. This is afforded by [FM94, Theorem 3]. (See also the proof of [EL23, Theorem 9.2.3] for further explanation.) Therefore, (5.6) restricts to an isomorphism
@@
The value of the constant $C^{2}$ in § 5.3 (which is also the value of the constant $C$ in Theorem 1.1.1 and Theorem 1.2.1) can be explicitly computed in terms of the point at which the homology of a certain Hurwitz space stabilizes. The computation of this constant does not use our work, but entirely follows from the proof of [EVW16]. For explicitly determining the constant, it was also useful for us to consult [RW20] and [EL23]. It may be possible to reduce the value of the constant $C$, so what we describe below is really an upper bound for $C^{2}$.
@@ A.4. A simple topological model
It is easy to see that for each $01>\epsilon>$ 0, $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ admits the structure of a CW-complex of dimension $<\frac{2}{\epsilon}$. This is because each component of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is a union of disks, and the gluing relations to form $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ are along subcomplexes. It is finite dimensional since $\epsilon$ puts a bound on the number of points in the configurations that are allowed. For a suitably chosen CW structure on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$, the inclusion $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is that of a subcomplex, showing the lemma. ∎

### v2

@@ 5.2. A preliminary lemma on point counting
In what follows, if $x$ is a complex number, we use $\|x\|$ to denote its absolute value, while if $X$ is a set, we use $|X|$ to denote its cardinality. For $\mathscr{X}$ a Deligne–Mumford stack, we follow the standard convention that $|\mathscr{X}(\mathbb{F}_{q})|$ denotes $\sum_{x\in\mathscr{X}(\mathbb{F}_{q})}\frac{1}{|\operatorname{Aut}(x)|}$, so points are weighted inversely proportionally to their automorphism groups. Of course, when $\mathscr{X}$ is a scheme, this agrees with the usual notion of $|\mathscr{X}(\mathbb{F}_{q})|$.
@@
Fix a prime power $q$ and a sequence $\{Y_{n}\}_{n\in S}$ of nonempty smooth Deligne–Mumford stacks over $\operatorname{Spec}\mathbb{F}_{q}$ with $Y_{n}$ of pure dimension $n$, where $n$ traverses over an infinite sequence of increasing integers $S\subset\mathbb{Z}_{\geq 0}$. Suppose that the following cohomological conditions are satisfied:
@@ 5.3. Applying the lemma to count points on Hurwitz spaces
Suppose $G$ is a finite group, $c\subset G$ is a conjugacy class closed under invertible powering, and $(G,c)$ is non-splitting. Let $q$ be an prime power satisfying $\gcd(q,|G|)=1$. If $Y_{n}\subset\operatorname{CHur}^{G,c}_{n,\mathbb{F}_{q}}$ is a union of connected components which has $r$ geometrically irreducible components for each $n$ sufficiently large, there exist constants $C,I,J$, all independent of $n$ and $q$, but possibly depending on $r$, so that, as long as $q>C^{2}$,
@@
once we exhibit a suitable normal crossings compactification of configuration space. This is afforded by [EL23, Corollary B.1.4]. Therefore, (5.6) restricts to an isomorphism
@@
The value of the constant $C^{2}$ in § 5.3 (which is also the value of the constant $C$ in Theorem 1.1.1 and Theorem 1.2.1) can be explicitly computed in terms of the point at which the homology of a certain Hurwitz space stabilizes. The computation of this constant does not use our work, but entirely follows from the proof of [EVW16]. For explicitly determining the constant, it was also useful for us to consult [RW20] and [EL23]. It may be possible to reduce the value of the constant $C^{2}$, so what we describe below is really an upper bound for $C^{2}$.
@@ A.4. A simple topological model
It is easy to see that for each $01>\epsilon>$ 0, $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ admits the structure of a CW-complex of dimension $<\frac{2}{\epsilon}$. This is because each component of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is a union of discs, and the gluing relations to form $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ are along subcomplexes. It is finite dimensional since $\epsilon$ puts a bound on the number of points in the configurations that are allowed. For a suitably chosen CW structure on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$, the inclusion $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is that of a subcomplex, showing the lemma. ∎